OpenStudy (anonymous):

simplify 4sqrt400/4sqrt5 Can someone walk me through the steps?

OpenStudy (anonymous):

\[\frac{ 4\sqrt{400} }{4\sqrt{5} }\]

OpenStudy (loser66):